OpenStudy (anonymous):

What is the focal length of a makeup mirror that produces a magnification of 1.50 when a person’s face is 12.0 cm away?

OpenStudy (shamim):

Its a concave mirror. Because concave mirror has magnitude more than 1. Convex mirror dont have magnification more than 1

OpenStudy (shamim):

Nd the image of a concave mirror is real and opposit. So magnification m=-1.5

OpenStudy (shamim):

But m=-v/u -v/u=-1.5 v=1.5*u v=1.5*12=18cm Here given u=18cm

OpenStudy (shamim):

U need to use the equation 1/v+1/u=1/f f=?

OpenStudy (shamim):

Response plz